CJ) r + 'Y' = (% +,x - %)' + y2>
.

Grade:12th pass

1 Answers

Pawan Prajapati
3 years ago
(i) Consider the identity (.r + y)3 = xs + ya + !hy(.r + y) ,rS + yS = (% + y)S _ 3x:y{.r + y) = (.r + y)((% + y)l - 3%)'1 .r8 + y3 = (% + y) t.r2 + y2 + 2xy - 3.ryJ = C% + :iX -%Y + y). (ii} Consider the identity (.r -y'fJ = .r3 -y3 - 3%)'(.r - y) .r3 -y3 = (.r - y '/3 + 3%)'(% - y) = (% -y)l(.r -y'f/- + 3%)'1 .r3 - y 8 = <.r -Y> tr + ,a - 2x:Y + :uy1 = (.r - y)(z2 + %)' + y2).